Can somebody please explain to me how to work this equation?
3/8 x-1/2=5/6-x

Answer by stanbon(75887)

work this equation?
3/8 x-1/2=5/6-x
----
Multiply thru by 24 to get:
3*3x - 12 + 5*4 - 24x
9x -12 = 20 - 24x
33x = 32
x = 32/33
-----------------
Cheers,
Stan H.